# PyraTrans: Attention-Enriched Pyramid Transformer for Malicious URL Detection

### v2

@@ Abstract
Although advancements in machine learning have driven the development of malicious URL detection technology, current techniques still face significant challenges in their capacity to generalize and their resilience against evolving threats. In this paper, we propose PyraTrans, a novel method that integrates pretrained Transformers with pyramid feature learning to detect malicious URL. PyraTrans utilizes a pretrained CharBERT as its foundation and is augmented with three interconnected feature modules: 1) Encoder Feature Extraction, extracting multi-order feature matrices from each CharBERT encoder layer; 2) Multi-Scale Feature Learning, capturing local contextual insights at various scales and aggregating information across encoder layers; and 3) Spatial Pyramid Attention, focusing on regional-level attention to emphasize areas rich in expressive information. The proposed approach addresses the limitations of the Transformer in local feature learning and regional relational awareness, which are vital for capturing URL-specific word patterns, character combinations, or structural anomalies. In several challenging experimental scenarios, the proposed method has shown significant improvements in accuracy, generalization, and robustness in malicious URL detection. For instance, it achieved a peak F1-score improvement of 40% in class-imbalanced scenarios, and exceeded the best baseline result by 14.13% in accuracy in adversarial attack scenarios. Additionally, we conduct a case study where our method accurately identifies all 30 active malicious web pages, whereas two pior SOTA methods miss 4 and 7 malicious web pages respectively. Codes and data are available at: https://github.com/Alixyvtte/PyraTrans.
Index Terms:
@@
Cybersecurity, Phishing Detection, Transformer, Multi-Scale Learning, Pyramid Attention
I Introduction
@@
Malicious URLs, systematically engineered by cybercriminals for illicit activities such as scams, phishing, spam, and malware distribution, constitute a significant cybersecurity risk. These URLs directly threaten user and organizational security, leading to privacy breaches, data theft, extortion, and compromising the integrity of devices and networks. Vade’s 2023-Q3 report highlights a substantial escalation in phishing and malware incidents. The period saw malware volumes nearing the highest quarterly record, second only to Q4 2016’s 126.8 million instances. A notable 173% increase in phishing incidents was reported over the previous quarter, totaling 493.2 million, the highest for any Q3 since systematic tracking began in 2015 [1].
@@
Predominantly, cyber attackers leverage deceptive hyperlinking as a key strategy in phishing schemes, often imitating credible entities like Microsoft, Google, and Facebook [2]. The ability to alter the display text of hyperlinks in HTML exacerbates the threat by camouflaging the true malicious nature of these URLs. This tactic poses a significant challenge to the detection of malicious URLs. Traditional database-centric detection platforms, such as Phishtank, reliant on community submissions, often face delays in identifying and mitigating new threats. The urgency for advanced systems capable of adapting to the dynamic cybersecurity environment and effectively identifying deceptive malicious URLs is highlighted by this lag.
@@
Traditional methods for detecting malicious URLs, such as blacklist, heuristic, and rule-based approaches, face limitations as they require a understanding of URL structures and nature (as depicted in Figure 1[3, 4]. Blacklist detection encounters issues such as delayed updates and struggles with identifying unknown malicious URLs. Heuristic methods show performance variations across different malicious URL samples, while rule-based detection needs manual maintenance, making it challenging to address emerging threats and their variations[5, 6]. These limitations highlight the need for utilizing more advanced machine learning methods. On the other hand, earlier studies illustrates that malicious URLs display highly discernible string patterns, including length, the quantity of dots, and specific words [7, 8]. These patterns play a vital role in threat analysis and provide the groundwork for training sophisticated classifiers.
@@
TABLE I: Example of the BERT token sequence extraction from amazon web page.
@@
Advancements in deep learning have significantly propelled the development of malicious URL detection systems [9]. Utilizing stacked nonlinear layers, deep learning [10, 11] learns to extract abstract, dense representations from raw initial feature vectors, such as character-level, word-level, or n-gram-based features, circumventing the need for intricate expert feature engineering. Convolutional Neural Networks (CNNs) have been the predominant deep learning models in prior malicious URL detection efforts like URLNet [12], TException [13], and Grambeddings [14], where CNNs capture local URL patterns to achieve SOTA performance. However, the technical limitations of CNNs make it increasingly difficult to achieve significant improvements in the development of CNN-based malicious URL detection models.
Recently, pretrained Transformer frameworks, such as BERT [15], have demonstrated promising prospects in various fields including natural language understanding, computer vision, and bioinformatics. This novel computational architecture and training paradigm offer new opportunities in malicious URL detection [16, 17]. They not only have stronger contextual learning capabilities but are also purely data-driven, end-to-end models. Table I presents an example of a BERT token sequence generated from a URL. However, despite the success of pretrained Transformer across multiple disciplines, their technological advantages have yet to be fully explored and manifested in malicious URL detection. Additionally, when applying Transformers to URLs, extra design considerations are essential for local feature learning and focused attention on key segments, as these aspects are crucial for capturing specific word patterns, character combinations, or structural anomalies in URLs.
@@
In this study, we propose PyraTrans, a novel appraoch developed on the foundation of pretrained Transformers, aimed at effectively detecting malicious URLs. This method exploits the various encoding layers of pre-trained Transformers and amalgamates them with multi-scale feature extraction and pyramid attention. Our approach, when tested across a broad range of public benchmarks datasets, exhibited superior performance over previous methods, proving to be more generalizable and robust, particularly in challenging situations like adversarial attacks, cross-dataset evaluation, class-imbalanced learning. We also conduct a case study to evaluate the model in real-world applications.
@@
  • Our method establishes new SOTA performance in malicious URL detection. Throughout extensive experimental evaluations, it consistently demonstrates superior generalization and robustness compared to previous methods. Additionally, it excels in learning under class imbalance and shows improved multi-classification performance, all while reducing the reliance on extensive training data.
  • This work introduces a straightforward multi-order URL feature generation approach by leveraging the multi-level encoding structure of Transformers. Additionally, we develop a multi-scale feature learning module to address the limitations of Transformers in extracting local string pattern from URLs.
  • We combine Transformers with spatial pyramid attention in our method, enabling it to assign regional-level attention. This allows our model to focus on segments with high information density in URLs while reducing the impact of noisy parts, such as random strings.
  • Our research reveals that even previously most popular methods, when subjected to comprehensive evaluations and practical case studies, can exhibit significant limitations. This underscores the critical need for robust case studies in assessing real-world applicability and for a more nuanced evaluation system that considers a method’s generalization and robustness.
The paper unfolds as follows: Section II conducts a literature review, while Section III outlines the datasets used. In Section IV, we provide a thorough explanation of the architecture and key components within our model. V details extensive experiments on malicious URL detection, benchmarking against baseline methods. A case study is then provided in Section V-E, and our conclusions are summarized in Section VII.
II Related Work
@@
II-A CNN-based Detection
Huang et al. [18] proposed a network that incorporates convolutional layers and two capsule network layers to learn the embedding representations of URLs. Wang et al. [19] combined CNNs and RNNs to extract key features for measuring content similarity, integrating these with static lexical features extracted from URLs using Word2Vec for their detection model. URLNet [12] introduced a dual-channel CNN approach for learning both character and word-level embeddings, combining these at the model’s top. This method not only achieved state-of-the-art performance at the time but also inspired numerous subsequent studies [13, 20, 21, 22], which all adopted the dual-channel feature extraction concept of URLNet. Recently, Bozkir et al. [14] developed Grambeddings, a neural network that effectively combines CNNs, LSTMs, and attention mechanisms. This network represents URL features through n-grams and has shown performance that surpasses URLNet in certain aspects.
@@ II-A CNN-based Detection
TABLE II: The statistical analysis of our datasets
@@
II-B Transformer-based Detection
@@
III Large Scale URL Dataset
The four datasets used for training, validation and testing are publicly available. These datasets share a similar schema, consisting of the browsing URL and a corresponding label indicating whether the URL has been identified as malicious or benign. Upon analyzing their statistical data, including sample size, average URL length, and top-level domain (TLD) types, as shown in TableII, we discovered certain variations in the URL data across these datasets, which contribute to a more comprehensive evaluation of our model.
GramBeddings Dataset: As provided by Grambeddings[14], this dataset comprises 800,000 samples, equally divided into 400,000 malicious and 400,000 benign URLs. Malicious URLs were collected from websites such as PhishTank and OpenPhish, spanning the period from May 2019 to June 2021. Long-Term and Periodical Sampling, as well as Similarity Reduction techniques, were applied to the malicious data. The benign URLs were iteratively crawled from Alexa and the top 20 most popular websites in 20 different countries, and then randomly sampled. As a result, this dataset presents the highest diversity and sample size compared to others, while also demonstrating an equal number of instances per class. As shown in the table, the average length of malicious URLs is significantly longer than that of benign URLs, approaching twice the length, ensuring the similarity of the class-level average length distribution. In terms of top-level domain (TLD) features, because different domains are generally more difficult to obtain in malicious URLs, this dataset has improved the domain-level diversity of malicious URLs by setting a low ratio of unique domains to total domains, achieving a ratio similar to that of benign URLs, with .com at 60.10% and ccTLDs at 11.82%, whereas for benign URLs, it is .com at 52.17% and ccTLDs at 12.04%.
@@ III Large Scale URL Dataset
The Kaggle 2 dataset consists of four classes: benign (428,079), defacement (95,306), phishing (94,086), and malicious (23,645). The benign class contains positive samples, while the other three classes contain negative samples of different types. We observe that the .com TLDs are dominant in benign URLs, accounting for 74.27% of the total. The ccTLDs are slightly more frequent in this dataset (6.61%) than in the other two datasets, while the other gTLDs represent 19.12% of the benign URLs. For the negative samples, the .com TLDs are less prevalent, with a frequency of 46.62% across all three classes. The ccTLDs and other gTLDs have higher frequencies of 7.74% and 45.65%, respectively, in the negative samples than in the benign ones.
The distinctive composition and Top-Level Domain (TLD) distribution in each dataset offer a comprehensive foundation for assessing the efficacy of our method across diverse web domains. This enables robust testing under different real-world scenarios.
IV Methodology
The proposed method employs the CharBERT network structure as the backbone network, integrating an encoder feature extraction module, a multi-scale learning module, and a Spatial Pyramid Attention module.
@@ IV Methodology
Figure 2: The architectural diagram of BiGRU.
Figure 3: PyraTrans: Composed of Four Core Components. CharBERT, the backbone network for learning character and subword-level features; Encoder Feature Extractor, Multi-Scale Feature Learning, and Spatial Pyramid Attention module for acquiring multi-order, multi-scale, attention-weighted features.
IV-A Backbone Network
@@
The character-aware embedding of each token is primarily generated through two components: the encoding of individual characters and subword units. These two components are integrated via a dual-channel architecture. To establish contextual character embeddings, we utilize a bidirectional Gated Recurrent Unit (BiGRU) layer. The BiGRU employs a bidirectional recurrent neural network with only the input and forget gates[29]. The architecture diagram of the BiGRU is depicted in Figure 2.
@@
The heterogeneous interaction module fuses and separates the token and character representations after each transformer layer. The structure shown in Fig4. This module uses different fully-connected layers to transform the representations, and then concatenates and integrates them by using a CNN layer, as follows:
@@
Next is a fully connected layer with GELU activation[30], used to restore the fused features to two channels, with residual connection added to preserve the original information of each channel.
@@
IV-B Encoder Feature Extraction
Pre-trained language models such as BERT use multiple layers of Transformer encoders to learn semantic knowledge from large-scale corpora, and then fine-tune them for specific downstream tasks. Most BERT-based classification models depend on the [CLS] feature of the final layer, which summarizes the semantic information of the whole input sequence. However, Jawahar et al. [31] show that BERT can learn various information across layers, such as phrase-level details in lower layers, syntactic information in middle layers, and rich semantic features in higher layers. They apply k-means clustering to the BERT layer representations and measure the cluster quality by using Normalized Mutual Information (NMI). They find that lower BERT layers are better at encoding phrase-level information, as indicated by higher NMI scores, as shown in Table III. Deeper encoder layers are more effective in handling long-range dependency information.
@@ IV-B Encoder Feature Extraction
TABLE III: Performance of span representation clustering derived from various layers of BERT[31]
Layer 1 2 3 4 5 6
NMI 0.38 0.37 0.35 0.3 0.24 0.2
Layer 7 8 9 10 11 12
NMI 0.19 0.16 0.17 0.18 0.16 0.19
@@
$k_{1}$ $=\begin{bmatrix}x_{11}^{1}&x_{12}^{1}&\cdots&x_{1d}^{1}\\ x_{21}^{1}&x_{22}^{1}&\cdots&x_{2d}^{1}\\ \vdots&\vdots&\ddots&\vdots\\ x_{w1}^{1}&x_{w2}^{1}&\cdots&x_{wd}^{1}\end{bmatrix},\quad u_{1}$ $=\begin{bmatrix}x_{11}^{2}&x_{12}^{2}&\cdots&x_{1d}^{2}\\ x_{21}^{2}&x_{22}^{2}&\cdots&x_{2d}^{2}\\ \vdots&\vdots&\ddots&\vdots\\ x_{w1}^{2}&x_{w2}^{2}&\cdots&x_{wd}^{2}\end{bmatrix}.$ (14)
@@
IV-C Multi-scale Learning

### v3

@@ Abstract
While machine learning progress is advancing the detection of malicious URLs, advanced Transformers applied to URLs face difficulties in extracting local information, character-level information, and structural relationships. To address these challenges, we propose a novel approach for malicious URL detection, named TransURL, that is implemented by co-training the character-aware Transformer with three feature modules—Multi-Layer Encoding, Multi-Scale Feature Learning, and Spatial Pyramid Attention. This special Transformer allows TransURL to extract embeddings that contain character-level information from URL token sequences, with three feature modules contributing to the fusion of multi-layer Transformer encodings and the capture of multi-scale local details and structural relationships. The proposed method is evaluated across several challenging scenarios, including class imbalance learning, multi-classification, cross-dataset testing, and adversarial sample attacks. The experimental results demonstrate a significant improvement compared to the best previous methods. For instance, it achieved a peak F1-score improvement of 40% in class-imbalanced scenarios, and exceeded the best baseline result by 14.13% in accuracy in adversarial attack scenarios. Additionally, we conduct a case study where our method accurately identifies all 30 active malicious web pages, whereas two pior SOTA methods miss 4 and 7 malicious web pages respectively. The codes and data are available at: https://github.com/Vul-det/TransURL/.
Keywords:
@@
Malicious URL Detection Multi-Scale Learning Transformer Pyramid Attention.
1 Introduction
@@
Malicious URLs, systematically engineered by cybercriminals for illicit activities such as scams, phishing, spam, and malware distribution, constitute a significant cybersecurity risk. These URLs directly threaten user and organizational security, leading to privacy breaches, data theft, extortion, and compromising the integrity of devices and networks. Vade’s 2023-Q3 report indicates a significant rise in phishing and malware, with malware volumes approaching a record high since Q4 2016, and phishing incidents increasing by 173% from the previous quarter, reaching 493.2 million, the highest Q3 since 2015 [10].
@@
In general, cybercriminals leverage deceptive hyperlinking as a key strategy in phishing schemes, often imitating credible entities such as Microsoft, Google, and Facebook [8]. The ability to alter the display text of hyperlinks in HTML exacerbates the threat by camouflaging the true malicious nature of these URLs. This tactic poses a significant challenge to the detection of malicious URLs.
@@
Traditional detection methods like Phishtank and blacklist, heuristic, and rule-based approaches face delays and limitations in identifying new threats, as they depend on known URL structures and manual updates, struggling with novel malicious URLs [29, 20, 24, 27]. These limitations highlight the necessity for advanced machine learning techniques in the ever-evolving cybersecurity landscape. On the other hand, earlier studies illustrates that malicious URLs display highly discernible string patterns(such as Fig. 1) , including length, the quantity of dots, and specific words [17, 1]. These patterns play a vital role in threat analysis and provide the groundwork for training sophisticated classifiers.
@@
Table 1: Example of the BERT token sequence extraction from amazon web page.
@@
The advancement of deep learning has significantly propelled the development of malicious URL detection systems [18, 25, 4, 26, 33], with Convolutional Neural Networks (CNNs) being the cornerstone in previous efforts, as exemplified by URLNet [19], TException [34], and GramBeddings [2], which remain among the SOTA methods for malicious URL detection. However, the inherent technical constraints of CNNs have increasingly made it challenging to achieve substantial improvements in CNN-based malicious URL detection models.
Recently, pretrained Transformer frameworks like BERT (Bidirectional Encoder Representations from Transformers)[7] have expanded their exceptional sequence modeling capabilities beyond natural language processing into various domains [28, 3, 31]. This innovative computational architecture and training paradigm offer enhanced contextual learning capabilities and function as purely data-driven end-to-end models (Table 1 provides an example of a BERT token sequence generated from a URL). However, the standard Transformer encounters specific challenges in malicious URL detection: 1) It struggles to capture character-level information due to its token-based input mechanism, critical for identifying subtle alterations in URLs. 2) Transformers are less effective than CNNs at detecting local patterns crucial for identifying potentially malicious substructures in URLs. 3) Transformers lack the capability to directly discern the hierarchical structure inherent in URLs.
@@
This paper introduces TransURL, addressing the challenges faced by Transformers in malicious URL detection tasks. TransURL, built on a specialized Transformer architecture, leverages embeddings from all encoding layers, integrating advanced multi-scale feature learning with spatial pyramid attention mechanisms to achieve state-of-the-art malicious URL detection.
@@
  • The proposed method achieves SOTA (State of the Art) performance across a range of challenging scenarios, including class imbalance, small sample learning, multi-classification, cross-dataset validation, and adversarial sample attacks, comparing previous best methods. Furthermore, its practicality is further demonstrated through case studies.
  • The method introduced operates on a character-perceptive Transformer structure, effectively deriving embeddings that contain both subword and character-level information from a token sequence within a URL, all without relying on manual dual-input configuration.
  • Our method is the first to dynamically fuse multiple encoding layers of a deep Transformer framework, achieving nuanced multi-level feature extraction from URL sequences, and it provides empirical proof of the performance improvements attributed to this fusion of information.
  • We propose a joint training framework that combines the Transformer with multi-scale convolution and spatial pyramid attention techniques, leveraging their respective advantages. This innovative framework is capable of long-distance sequence modeling, supporting advanced multi-scale local feature extraction and global information aggregation.
  • Our rigorous experimental setup exposes that even the previously most effective methods have their limitations in certain scenarios, providing a vital testing framework for constructing models designed for practical use.
The paper unfolds as follows: Section 2 conducts a literature review, while Section 3 outlines the datasets used. In Section 4, we provide a thorough explanation of the architecture and key components within our model. 5 details extensive experiments on malicious URL detection, benchmarking against baseline methods. A case study is then provided in Section 5.5, and our conclusions are summarized in Section 7.
2 Related Work
@@
2.1 CNN-based Detection
Huang et al. [14] proposed a network that incorporates convolutional layers and two capsule network layers to learn the embedding representations of URLs. Wang et al. [36] combined CNNs and RNNs to extract key features for measuring content similarity, integrating these with static lexical features extracted from URLs using Word2Vec for their detection model. URLNet [19] introduced a dual-channel CNN approach for learning both character and word-level embeddings, combining these at the model’s top. This method not only achieved state-of-the-art performance at the time but also inspired numerous subsequent studies [34, 35, 15, 40], which all adopted the dual-channel feature extraction concept of URLNet. Recently, Bozkir et al. [2] developed GramBeddings, a neural network that effectively combines CNNs, LSTMs (Long Short-Term Memory), and attention mechanisms. This network represents URL features through n-grams and has shown performance that surpasses URLNet in certain aspects.
@@ 2.1 CNN-based Detection
Table 2: The statistical analysis of our datasets.
@@
2.2 Transformer-based Detection
@@
Compared to pre-trained models such as BERT and RoBERTa, which also use transformer architectures, TransURL exhibits superior performance in capturing character-level information and local patterns crucial for malicious URL detection. This advantage stems from its dual-channel architecture, addressing the limitations of pre-trained models that rely solely on token-based input mechanisms. Additionally, TransURL incorporates multi-scale pyramid features, enabling the analysis of URLs at various granular levels to detect patterns and anomalies indicative of malicious activity. Traditional pre-trained models generally lack such mechanisms for multi-scale analysis, which can constrain their effectiveness. Moreover, while pre-trained models employ general attention mechanisms effective for text comprehension and generation, these mechanisms are not optimized for the unique structural characteristics of URLs. TransURL’s specialized attention mechanisms enhance its ability to discern intricate patterns and structural relationships within URL sequences, thereby improving detection performance.
3 Large Scale URL Dataset
The four datasets used for training, validation and testing are publicly available. These datasets share a similar schema, consisting of the browsing URL and a corresponding label indicating whether the URL has been identified as malicious or benign. Upon analyzing their statistical data, including sample size, average URL length, and top-level domain (TLD) types, as shown in Table 2, we discovered certain variations in the URL data across these datasets, which contribute to a more comprehensive evaluation of our model.
GramBeddings Dataset: As provided by GramBeddings[2], this dataset comprises 800,000 samples, equally divided into 400,000 malicious and 400,000 benign URLs. Malicious URLs were collected from websites such as PhishTank and OpenPhish, spanning the period from May 2019 to June 2021. Long-Term and Periodical Sampling, as well as Similarity Reduction techniques, were applied to the malicious data. The benign URLs were iteratively crawled from Alexa and the top 20 most popular websites in 20 different countries, and then randomly sampled. As a result, this dataset presents the highest diversity and sample size compared to others, while also demonstrating an equal number of instances per class. As shown in Table 2, the average length of malicious URLs is significantly longer than that of benign URLs, approaching twice the length, ensuring the similarity of the class-level average length distribution. In terms of top-level domain (TLD) features, because different domains are generally more difficult to obtain in malicious URLs, this dataset has improved the domain-level diversity of malicious URLs by setting a low ratio of unique domains to total domains, achieving a ratio similar to that of benign URLs, with .com at 60.10% and ccTLDs (country code top-level domain) at 11.82%, whereas for benign URLs, it is .com at 52.17% and ccTLDs at 12.04%.
@@ 3 Large Scale URL Dataset
The Kaggle 2 dataset consists of four classes: benign (428,079), defacement (95,306), phishing (94,086), and malicious (23,645). The benign class contains positive samples, while the other three classes contain negative samples of different types. We observe that the .com TLDs are dominant in benign URLs, accounting for 74.27% of the total. The ccTLDs are slightly more frequent in this dataset (6.61%) than in the other two datasets, while the other gTLDs (generic top-level domain) represent 19.12% of the benign URLs. For the negative samples, the .com TLDs are less prevalent, with a frequency of 46.62% across all three classes. The ccTLDs and other gTLDs have higher frequencies of 7.74% and 45.65%, respectively, in the negative samples than in the benign ones.
The distinctive composition and TLD distribution in each dataset offer a comprehensive foundation for assessing the efficacy of our method across diverse web domains. This enables robust testing under different real-world scenarios.
4 Methodology
The proposed method employs the CharBERT (Character-aware Pre-trained Language Model) network structure as the backbone network, integrating an encoder feature extraction module, a multi-scale learning module, and a Spatial Pyramid Attention module.
@@ 4 Methodology
Figure 2: A Network Structure Diagram of the BiGRU Module.
Figure 3: TransURL: Composed of Four Core Components. CharBERT, the backbone network for learning character and subword-level features; Encoder Feature Extractor, Multi-Scale Feature Learning, and Spatial Pyramid Attention module for acquiring multi-order, multi-scale, attention-weighted features.
4.1 Backbone Network
@@
The character-aware embedding of each token is primarily generated through two components: the encoding of individual characters and subword units. These two components are integrated via a dual-channel architecture. To establish contextual character embeddings, we utilize a bidirectional Gated Recurrent Unit (BiGRU) layer. The BiGRU employs a bidirectional recurrent neural network with only the input and forget gates[6]. The architecture diagram of the BiGRU is depicted in Fig. 2.
@@
The heterogeneous interaction module fuses and separates the token and character representations after each transformer layer. The structure shown in Fig. 4. This module uses different fully-connected layers to transform the representations, and then concatenates and integrates them by using a CNN layer, as follows:
@@
Next is a fully connected layer with GELU activation [13], used to map the fused features onto two channels. A residual connection is added to preserve the original information of each channel.
@@
4.2 Encoder Feature Extraction
Pre-trained language models such as BERT use multiple layers of Transformer encoders to learn semantic knowledge from large-scale corpora, and then fine-tune them for specific downstream tasks. Most BERT-based classification models depend on the [CLS] feature of the final layer, which summarizes the semantic information of the whole input sequence. However, Jawahar et al. [16] show that BERT can learn various information across layers, such as phrase-level details in lower layers, syntactic information in middle layers, and rich semantic features in higher layers. They apply k-means clustering to the BERT layer representations and measure the cluster quality by using Normalized Mutual Information (NMI). They find that lower BERT layers are better at encoding phrase-level information, as indicated by higher NMI scores[16], as shown in Table 3. Deeper encoder layers are more effective in handling long-range dependency information.
@@ 4.2 Encoder Feature Extraction
Table 3: Performance of span representation clustering derived from various layers of CharBERT.
Layer 1 2 3 4 5 6 7 8 9 10 11 12
NMI 0.38 0.37 0.35 0.3 0.24 0.2 0.19 0.16 0.17 0.18 0.16 0.19
@@
$k_{1}$ $=\begin{bmatrix}x_{11}^{1}&x_{12}^{1}&\cdots&x_{1d}^{1}\\ x_{21}^{1}&x_{22}^{1}&\cdots&x_{2d}^{1}\\ \vdots&\vdots&\ddots&\vdots\\ x_{w1}^{1}&x_{w2}^{1}&\cdots&x_{wd}^{1}\end{bmatrix},\quad u_{1}$ $=\begin{bmatrix}x_{11}^{2}&x_{12}^{2}&\cdots&x_{1d}^{2}\\ x_{21}^{2}&x_{22}^{2}&\cdots&x_{2d}^{2}\\ \vdots&\vdots&\ddots&\vdots\\ x_{w1}^{2}&x_{w2}^{2}&\cdots&x_{wd}^{2}\end{bmatrix}$ (14)
@@
4.3 Multi-scale Learning